A 50kg cat statue falls from a second floor (10m high) window. With what momentum does it impact the ground with?


1
Expert's answer
2021-03-16T11:39:20-0400

Let's first find the impact velocity of the cat statue:


"PE=KE,""mgh=\dfrac{1}{2}mv^2,""v=\sqrt{2gh}=\sqrt{2\cdot9.8\ \dfrac{m}{s^2}\cdot10\ m}=14\ \dfrac{m}{s}."

Finally, we can find the momentum of the cat statue:


"p=mv=50\ kg\cdot14\ \dfrac{m}{s}=700\ \dfrac{kgm}{s}."